Do not round intermediate calculations. Give your final answer(s) to three significant figures. Knowing that beam A B is an S200 x 34 rolled shape and that P = 40 kN, L = 4 m, and E-200 GPa, determine the slope at A x 10rad
Answer #1

Given is:-

P = 40 × 103 N. L = 4m. E-200 103Mpa

The exaggerated shape of the elastic curve for the beam is as shown below:

L-2 0 dy盂 L-2

Now,

Raz-

M_x = rac{P}{2}x (at x = 0, L/2)

Applying deflection equation we get

EI rac{d^2y}{dx^2} = Mx

dy P d2

by integrating we get

EI rac{dy}{dx} = rac{P}{2} imes rac{x^2}{2} + C_1

d.x 4 + C1

at x = L/2 dy/dx =0

we get

Eldy=0

PL2 16

or

PL2 16 C1 =-

now

at x=0   heta = heta_{max} = rac{PL^2}{16EI} eq-1

by obtaining the value of moment of inertia of S200*34 roller shape from American standard shapes table we get

I-26.9 x 10mm

by plugging all the values in eq-1 we get

(40 x 10) (4 x 1000)2 (16) (200 x 103) (26.9 x 106)

which gives us

θΑ 7.434 × 10-rad! clockwise